OpenStudy (anonymous):

x^3-2x^2-3x+6 Can you show me the steps on how to find the zeros.

sam (.sam.):

Did you use trial by error before?

OpenStudy (anonymous):

no

Directrix (directrix):

x^3-2x^2-3x+6 = x² ( x - 2 ) - 3 ( x - 2 ) Note that (x-2) is a common factor.

Directrix (directrix):

x² ( x - 2 ) - 3 ( x - 2 ) = (x - 2) * ( x² - 3)

Directrix (directrix):

(x - 2) * ( x² - 3) = 0 x - 2 = 0 OR ( x² - 3) = 0 Solve each of the two above equations for x.